OpenStudy (anonymous):

A placekicker kicks a football upward at an angle of 20 degrees and a speed of 5.0 m/s. At what other angle will he have to kick a second ball with the same speed to reach the same distance downrange? Ignore air resistance! 30 degrees, 45 degrees, 60 degrees, or 70 degrees **not sure how to solve! thanks! :)

OpenStudy (mrnood):

OK - first you need to find out how long th eball is in their for th efirst kick

OpenStudy (mrnood):

Veritically the initial speed is 5 sin 20 and the time is given by v=u+at v=0 at peak height and a = g this gives the time to peak: total time is double that time

OpenStudy (michele_laino):

please you have to use the formula for the range, namely \[range=\frac{ v _{0}^{2} }{ g }\sin (2 \alpha)\] where g is the gravity, v_0 is the initial speed and alpha is tha angle of launch. So please calculate the range for your first football

OpenStudy (anonymous):

ah okay wait so what am i plugging in?

OpenStudy (michele_laino):

you have to use the formula above, and please insert your numerical data, namely: Alpha=20, v_0=5, and g=9.81

OpenStudy (anonymous):

oh okay so i get 5^2/9.81*sin(2*20)=1.89886?

OpenStudy (michele_laino):

I got 1.639

OpenStudy (anonymous):

ohh okay! so what happens from there?

OpenStudy (michele_laino):

Now in order to find your answer you have to solve this equation: \[\frac{ 5^{2} }{ 9.81 }\sin (2 x)=1.639\] please try to solve that equation for x

OpenStudy (anonymous):

okay:) so i get to here, is this correct? sin(2x)=0.643 ?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

and then i forget how to isolate the x here:(

OpenStudy (michele_laino):

you have to to this: \[x=\frac{ 1 }{ 2 }\sin^{-1} (0.643)\]

OpenStudy (anonymous):

ohh okay:) and then now solve and get 0.349?

OpenStudy (michele_laino):

please you shoul get x=20

OpenStudy (anonymous):

oh! yes sorry, my calc was in radian mode! fixed it:) what happens from here now?

OpenStudy (michele_laino):

please note that, I cna write this: \[1.639=\frac{ 25 }{ 9.81 }2*\sin 20 \cos 20\] do you agree?

OpenStudy (michele_laino):

oops... I can write...

OpenStudy (michele_laino):

it is the formula we have already used

OpenStudy (anonymous):

okay:)

OpenStudy (michele_laino):

now, please note that: \[\cos 20=\cos(90-70)=\cos 90 \cos 70-\sin 90 \sin 70= \sin 70\] and: \[\sin 20=\sin (90-70)=\sin 90 \cos 70-\cos 90 \sin 70=\cos 70\] so, substituting into the formula for range, we have: \[range=\frac{ 25 }{ 9.81 }*2 \sin 20 \cos 20=\] \[=\frac{ 25 }{ 9.81 }2*\cos 70 *\sin 70=\frac{ 25 }{ 9.81 }*2 \sin 70 \cos 70\] so other angle is 70°

OpenStudy (anonymous):

ahh okay:) thank you @Michele_Laino :)

OpenStudy (michele_laino):

thank you!! :) @iheartfood

OpenStudy (anonymous):

:)